OpenStudy (anonymous):

L'Hopital's Rule Problem: lim(x->0) (1/((x^(1/5))(lnx))) I know the answer is -infinity, but I can't figure out how to get there.

OpenStudy (anonymous):

\[\lim_{x\to 0}\frac{1}{\sqrt[5]{x}\ln(x)}\]?

OpenStudy (anonymous):

Yes, sorry if it was unclear..

OpenStudy (anonymous):

and it is approaching 0^(+)

OpenStudy (anonymous):

ok well you cannot use l'hopital as it is , since it is not in the form \(\frac{0}{0}\) or \(\frac{\infty}{\infty}\) we will have to rewrite it some how to make it one of those

OpenStudy (anonymous):

usual gimmick is to take the reciprocal of one of those things i would pick rewriting as \[\lim_{x\to 0}\frac{1}{\sqrt[5]{x}\ln(x)}=\lim_{x\to 0}\frac{\frac{1}{\sqrt[5]{x}}}{{\ln(x)}}\]

OpenStudy (anonymous):

I wrote it as (1/x^(1/5))/ln(x)... basically what you did..

OpenStudy (anonymous):

now it looks like \(\frac{-\infty}{-\infty}\) and you can use l'hopital

OpenStudy (anonymous):

oh ok then you should be in good shape

OpenStudy (anonymous):

so that's what I got, but apparently the answer should be -infinity.. not positive infinity.

OpenStudy (anonymous):

hmm

OpenStudy (anonymous):

ok i'm going to write out what I worked out.. and please tell me if it's correct.. i'll use that equation drawing thing

OpenStudy (anonymous):

how come i am getting zero? am i doing something wrong?

OpenStudy (anonymous):

oh yeah i am

OpenStudy (anonymous):

ok this drawing is taking way too long.. I got: (d/dx)(1/x^(1/5)) = -1/(5*(x^(6/5)) then d/dx of lnx is 1/x

OpenStudy (anonymous):

working that out I got 0 at first, then I made some random adjustments to the order (putting ln(x)) in the numerator and ended up with positive infinty

OpenStudy (anonymous):

yeah your exponent is wrong, it should be negative

OpenStudy (anonymous):

it is.. i wrote -1/... making it negative

OpenStudy (anonymous):

\[\large \frac{1}{x^{\frac{1}{5}}}=x^{-\frac{1}{5}}\] the derivative is \[\large -\frac{1}{5}x^{-\frac{6}{5}}\]

OpenStudy (anonymous):

you get \[\frac{-\frac{1}{5}x^{-\frac{6}{5}}}{\frac{1}{x}}\]

OpenStudy (anonymous):

clean it up and you get \[-\frac{1}{5x^{\frac{1}{5}}}\]

OpenStudy (anonymous):

so if you have -(1/0) that gives you -infin?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

denominator is going to zero through positive numbers, since you are taking \[\lim_{x\to 0^+}\]

OpenStudy (anonymous):

and you have a minus sign out front, making the whole thing negative

OpenStudy (anonymous):

so i just spent the past 30 minutes redoing this problem because i didn't know that.. i feel.. inadequate.. thank you for the help, though!!

OpenStudy (anonymous):

lol yw